OpenStudy (anonymous):

A conic section has the equation x2 + y2 + 12x + 8y = 48. Determine the following: type of conic, domain and range, axes of symmetry, and center. Show your work.

OpenStudy (anonymous):

So I know it is a circle with the equation (x+6)^2+(y+4)^2=10^2 I'm pretty sure that it has infinite axis of symmetry since it is a circle, right? How do I find domain and range and center?

OpenStudy (anonymous):

@Mertsj

OpenStudy (mertsj):

Write it in this form: \[(x-h)^2+(y-k)^2=r^2\]

OpenStudy (anonymous):

Ya I did (x+6)^2+(y+4)^2=10^2 :P

OpenStudy (mertsj):

The center of that circle is (h,k) and the radius is r

OpenStudy (anonymous):

So center is (6,4)

OpenStudy (mertsj):

(x-(-6))^2+(y-(-4))^2=10^2

OpenStudy (anonymous):

Ohh I see, (-6,-4)

OpenStudy (mertsj):

yes

OpenStudy (anonymous):

Domain and range? and axis of symmetry?

OpenStudy (mertsj):

Now the domain: If the center is (-6,-4) and the radius is 10, what x values would be on the circle?

OpenStudy (anonymous):

-16 to 4?

OpenStudy (mertsj):

Exactly.

OpenStudy (anonymous):

So domain is -16<x<4?

OpenStudy (mertsj):

So domain is [-16.4] in interval notation

OpenStudy (mertsj):

yes

OpenStudy (mertsj):

And apply the same thought process for the range.

OpenStudy (anonymous):

And range -14<y<6?

OpenStudy (mertsj):

yes

OpenStudy (anonymous):

And just to check, for the axis of symmetry, I would but infinite?

OpenStudy (mertsj):

Does it want to know how many axes of symmetry?

OpenStudy (anonymous):

No, it wants to know what the axis of symmetry is

OpenStudy (mertsj):

That is a weird question. Any line through the center is an axis of symmetry

OpenStudy (mertsj):

So do you have answer choices?

OpenStudy (anonymous):

I think I'll just put "all lines through center axis".. and no, no choices on this one

OpenStudy (mertsj):

That sounds right. You might say, "Any line that contains (-6,-4) is an axis of symmetry and so there is an infinite number of axes of symmetry."

OpenStudy (anonymous):

Thanks for the 100th time haha, you're an amazing helper!

OpenStudy (mertsj):

yw, my pleasure to work with such a smart fellow.